OpenStudy (anonymous):

Plz help for a medal: An expression is shown below: f(x) = -16x2 + 22x + 3 Part A: What are the x-intercepts of the graph of the f(x)? Show your work. (2 points)

OpenStudy (caozeyuan):

you mean -16x^2?

OpenStudy (anonymous):

yea

OpenStudy (caozeyuan):

and x-intercept means y=0, so -16x^2+22x+3=0

OpenStudy (anonymous):

yes

OpenStudy (caozeyuan):

because 3 is prime we have (ax+3)(bx+1)=-16x^2+22x+3

OpenStudy (anonymous):

how did you get that?

OpenStudy (caozeyuan):

assume (ax+c))(bx+d)=-16x^2+22x+3

OpenStudy (anonymous):

oh ok but why did you put 3 and 1 into the slots?

OpenStudy (caozeyuan):

now, obvious cd=3, but 3 is prime which means the only possible answer is c=3,d=1, the other way round is the same, not a different answer

OpenStudy (anonymous):

oh so when it says "What are the x-intercepts of the graph" they would be c=3,d=1?

OpenStudy (caozeyuan):

we are not there yet, bear with me for a second

OpenStudy (anonymous):

ok

OpenStudy (caozeyuan):

now expand LHS we have ab=-16,3b+a=22

OpenStudy (caozeyuan):

we can recoginze that 8 and -2 is plausible answer

OpenStudy (anonymous):

why cause they both go into 16?

OpenStudy (caozeyuan):

since 3b+a is 22, a is -2, b is8

OpenStudy (caozeyuan):

sorry what is your question?

OpenStudy (anonymous):

Part A: What are the x-intercepts of the graph of the f(x)? Show your work.

OpenStudy (caozeyuan):

now since y is 0, either -2x+3 is 0 or 8x+1 is 0

OpenStudy (caozeyuan):

I dont understand "why cause they both go into 16?"

OpenStudy (anonymous):

I said that because you said 8 and 2 are possible anwser

OpenStudy (anonymous):

and 8 and 2 both go into 16

OpenStudy (caozeyuan):

ab=-16, so you want one+ and one -

OpenStudy (anonymous):

yea I get that part

OpenStudy (caozeyuan):

8 and 2 is not possible since it would give the coeffecient before x^2 16,not -16

OpenStudy (caozeyuan):

since 3b+a=22, we want to have 24-2, so a=-2 and b is 8

OpenStudy (anonymous):

I forgot it was a negative ha

OpenStudy (caozeyuan):

So now you should have factorized the quadratic equation

OpenStudy (caozeyuan):

since y=0, either the first factor is 0 or the second, set them each to 0 and solve for x, thats your x intercept

OpenStudy (anonymous):

oh ok so than what would be the anwser?

OpenStudy (caozeyuan):

you can calculate it yourself, it is simple once you factored the quadratic eqaution

OpenStudy (anonymous):

ok I got it now thanks and do you think you can help with part B and C?

OpenStudy (caozeyuan):

sure

OpenStudy (anonymous):

Part B: Is the vertex of the graph of f(x) going to be a maximum or minimum? What are the coordinates of the vertex? Justify your answers and show your work. (3 points) Part C: What are the steps you would use to graph f(x)? Justify that you can use the answers obtained in Part A and Part B to draw the graph. (5 points) @caozeyuan

OpenStudy (caozeyuan):

maxmium because secondary derivative is negative

OpenStudy (caozeyuan):

d^2y/dx^2=-32

OpenStudy (caozeyuan):

to find coord, you do first derivtive, set dy/dx=0, find, plug back in to find y

OpenStudy (caozeyuan):

I am not sure if you finished calc I or not

OpenStudy (caozeyuan):

well, a parabola is defined by 3 non collinear points, the two intercepts and the vertex are the three points, so yes it can be graphed

OpenStudy (anonymous):

could you just help me with part C and ill be good

OpenStudy (caozeyuan):

I am not sure about the standad steps, but I would draw a vertical line down to x axis from the vertex,this is the symmetry line of parabola, then connncet left intercept with vertex and right interest with vertex, make sure two curves are smooth and symmetrical

OpenStudy (anonymous):

Well you did great thanks for helping me out :)